OpenStudy (jocidiamandis):

i have a car wheel project and i need the measures of three car wheels but I can't seem to find the measures of three car wheels and once I do that do I just use the circumference to find the distance? im sorry im just really confused

OpenStudy (photon336):

well, I guess you would do it this way  once you get the radius you can easily find the circumference of the wheel. \[C = \pi*2r \]

OpenStudy (photon336):

@JociDiamandis

OpenStudy (jocidiamandis):

i see, how do i find out how many rotations it goes will it just be the circumference?

OpenStudy (photon336):

well, so my guess is that the number of rotations is the time it takes for the wheel to move a distance of the circumference. so I guess this would depend on how fast the wheel is moving.

OpenStudy (jocidiamandis):

ah okay because im trying to see how many rotations a wheel would go with certain measurements for 10,000 miles

OpenStudy (welshfella):

suppose we had a wheel with radius 1 foot then the circumference is 2 * pi *1 = 2pi feet

OpenStudy (welshfella):

so number of revolutions of the wheel = distance travelled / circumference Now 10,000 miles = 5280 * 10,000 feet so the answer will be 52,800,000 / 2 pi

OpenStudy (welshfella):

so grab your calculator and work it out Tell me what you get.

OpenStudy (jocidiamandis):

16,806,761.9?

OpenStudy (welshfella):

to get the correct answer note that you need the 2 measurements to be in the same unit. 9 in this case feet).

OpenStudy (jocidiamandis):

okay i see

OpenStudy (welshfella):

thats not correct - didid you miss out the 2 when you divided/

OpenStudy (welshfella):

i think you divided by pi instead of 2pi

OpenStudy (jocidiamandis):

would it be 8403380.9?

OpenStudy (welshfella):

right

OpenStudy (jocidiamandis):

okay thank you i understand now!

OpenStudy (welshfella):

good yw